Bonner company paid $855,000 for a tract of land which included a warehouse and office building. The land, warehouse and office building originally cost $280,000, $180,000, and $340,000 respectively and the current assessed values are $300,000, $200,000, and $400,000 respectively. How should Bonner calculate what value to record for Land?
A
300,000 / 900,000 x 855,000
B

280,000 / 800,000 x 855,000

C
300,000 / 800,000 x 855,000
D
280,000 / 900,000 x 855,000
Cost of goods available for sale was $900,000 and the gross profit rate was 20%. If sales were $800,000, what was ending inventory?
A

$0

B

$260,000

C

$160,000

D

$180,000
Answer #1

1. Total value = 300,000+200,000+400,000 = 900,000

Value to be recorded for land

= (300,000/900,000)*855,000

Option A is the answer

2.

Cost of Goods Sold = Sales - Gross profit = 800,000-20% = 640,000

Ending inventory = Cost of Goods available - Cost of Goods Sold

= 900,000 - 640,000

= 260,000

Option B is the answer